OpenStudy (anonymous):

what is the value of X in the equation -2x + 8 = _ -16

OpenStudy (anonymous):

Will give medals to those that help me.

OpenStudy (anonymous):

What is the value of x 6(-3 - x ) -2x = 14 wil give madals

OpenStudy (espex):

When you solved for 'x', what did you get?

OpenStudy (anonymous):

when I got to X i got 8 for the first one but you cant get 0 for X because then the equation doesn't make sense

OpenStudy (espex):

\(-2x + 8 = -16 \rightarrow -2x = (-16-8) \rightarrow -2x = -24 \rightarrow x = \frac{-24}{-2}\)

OpenStudy (espex):

What confused you on the second one?

OpenStudy (anonymous):

okay for the second one am i supposed to leave the 6 once i times it to -3? or would the equation turn into -18*6x-2x=14?

OpenStudy (espex):

With this as the equation, you first want to distribute the 6 into the parenthesis,6(-3 - x ) -2x = 14 \(6(-3-x) \rightarrow (6*(-3) + 6*(-x))\)

OpenStudy (anonymous):

okay after i do that would i combine the like terms so 6x and 2x

OpenStudy (espex):

Yes, exactly.

OpenStudy (espex):

So after distributing you would have \(-18 -6x-2x=14\)

OpenStudy (anonymous):

so after combining them i would have -18-8x=14

OpenStudy (espex):

Exactly, nicely done. :)

OpenStudy (anonymous):

okay now im lost again so i took -18-8x=14 and i added 8 to 14 so X would be left alone but what would i do after that?

OpenStudy (espex):

You cannot add the -8 and the -18 together since the -8 is multiplied by 'x'. So first you need to move the -18.

OpenStudy (anonymous):

okay so i didnt add the 8x to the -18 i added it to the 14

OpenStudy (anonymous):

so i did -18-8x=14 +8=+8 witch have me 22

OpenStudy (espex):

When you move the -18 from the left to the right, you have to 'add 18' to both sides. \(-18-8x=14 \rightarrow (+18) -18-8x=14 (+18) \rightarrow -8x = 14+18\)

OpenStudy (anonymous):

okay so now i have 8x=32

OpenStudy (anonymous):

so would X=4?

OpenStudy (espex):

Now since your -8 is multiplied by x, the only want to reverse multiplication is through division. That is the right equation though you dropped a sign. So now you must DIVIDE by a -8 on both sides. \(-8x=32 \rightarrow \frac{-8x}{-8}=\frac{32}{-8}\)

OpenStudy (espex):

You dropped the sign, but yes, excellent work. :)

OpenStudy (anonymous):

so its -4

OpenStudy (espex):

Yes. Good job.

OpenStudy (anonymous):

would you mined helping me with one more?

OpenStudy (espex):

Sure, what's the problem?

OpenStudy (anonymous):

X+5=-6x-16

OpenStudy (espex):

With this one you want to first get all the terms with an 'x' on one side and all the whole numbers onto the other.

OpenStudy (anonymous):

okay I dont how i would do that

OpenStudy (espex):

When moving from RHS (right hand side) to LHS (left hand side) of an equal sign you want to try and do one of two things. Either you add/subtract to both sides to get 0 or you multiply/divide to both sides to get 1.

OpenStudy (espex):

In this case you are only adding/subtracting to make your first move. So how would you get a -6x from the RHS to the LHS and a +5 from the LHS to the RHS?

OpenStudy (anonymous):

so would i subtract -6x to get to the LHS and then to get 5 to the RHS would i add?

OpenStudy (anonymous):

I am really lost on this one im not understanding it at all

OpenStudy (espex):

Remember that you're trying to get a '0' when adding/subtracting. So -6x -6x = -12x, which won't work because it changes the equation.

OpenStudy (espex):

\(X+5=-6x-16 \rightarrow (6x)+x+5=-6x-16=(6x)\) You see how we add 6x to each side of the equation? We "cancel" out the -6x on the right but we add it back in on the left.

OpenStudy (espex):

That should read: \(X+5=-6x-16 \rightarrow (6x)+x+5=-6x-16+(6x)\)

OpenStudy (anonymous):

okay so right now my equation should be 6x+5=16?

OpenStudy (espex):

It should be \(6x+x+5=16\)

OpenStudy (espex):

err, dropped a sign again. \(6x+x+5=-16\)

OpenStudy (espex):

So now can you get the 5 from the LHS to the RHS?

OpenStudy (anonymous):

so i would subtract 5 and move it to the right?

OpenStudy (espex):

You would subtract 5 from BOTH sides, yes.

OpenStudy (anonymous):

oh and would the -6x stay negative once it was moved from right to left?

OpenStudy (espex):

The 6x would not be negative because you are having to add 6x to cancel out that -6x from the RHS.

OpenStudy (anonymous):

okay and now would my equation be 6x+x=-16-5?

OpenStudy (espex):

Yes, exactly.

OpenStudy (anonymous):

okay what would my next step be?

OpenStudy (espex):

Next you need to consolidate terms. Add the 6x and x, then add the -16 and -5.

OpenStudy (espex):

Once that is done, you will have a number "multiplied" by x. The only way to reverse multiplication is through division....

OpenStudy (anonymous):

okay so now i should have 6x=-21?

OpenStudy (anonymous):

what would give me an answer of -3.5

OpenStudy (espex):

Remember, when you have a variable by itself, like 'x', that it is really the same as having 1x. So you had 6x+1x, what does that equal?

OpenStudy (anonymous):

7x

OpenStudy (espex):

Yes, and then x = ?

OpenStudy (anonymous):

so that would change my answer to -3

OpenStudy (espex):

Precisely. Excellent work. :)

OpenStudy (anonymous):

okay im sorry but i dont know how to do the ones with fractions could you help me with that one?

OpenStudy (espex):

Fractions?